OpenStudy (anonymous):

Which proportion would you use to find 38% of 90?

OpenStudy (anonymous):

How about \[\frac{38}{100} = \frac{x}{90}\]

OpenStudy (anonymous):

u suree?

OpenStudy (anonymous):

Yes. If you actually solve for x, 100x = 90(38) 100x = 3420 x = 34.2 which is 38% of 90